OpenStudy (anonymous):

solve 8^(y - 3) = 2^y

OpenStudy (ledah):

polynomials again? :D

OpenStudy (anonymous):

yes :P

OpenStudy (ledah):

would you like the real or alternative solution? :3

OpenStudy (anonymous):

the answer haha i'm so lost...

OpenStudy (anonymous):

\[8^{(y-3)}= 2^{y}\] the idea is to have the same base: \[2^{3(y-3)} =2^{y}\] 3(y-3) = y 3y-9 = y solve for y

OpenStudy (anonymous):

(8^y)(8^-3)=2^y 8^y * -24 =2^y -8^y =-8^y -24= -6^y -24/-6 4=y

OpenStudy (anonymous):

thank you !!!!!!!!!!!!! :D

OpenStudy (ledah):

what I got was this from wolfram

OpenStudy (anonymous):

np .

OpenStudy (anonymous):

8^y)(8^-3)=2^y

OpenStudy (anonymous):

8^y * -24 =2^y

OpenStudy (anonymous):

mulla you incorrect

OpenStudy (ledah):

I think the answer is 

OpenStudy (anonymous):

ledah you are correct, mull and aly check my methodology above and follow it

OpenStudy (anonymous):

if 9/2 is 4.5 then i got it right :P

OpenStudy (anonymous):

oohh am sorry..iits my bad! :(